OpenStudy (anonymous):

The function f(x) = 25(4)x represents the growth of a bird population every year in a national park. Kyle wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Kyle's purposes? f(x) = 25(4)the x over 2 power f(x) = twenty five halves(4)x f(x) = 25(4 to the one half power)2x f(x) = 50(2)x

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aum):

\[ f(x) = 25(4)^x ~~ \text{where x is in years. }\\ \text{But now it is changed so x represents half-year. }\\ \text{So x has to be divided by 2 to convert it to a year in the original formula:} \\ \large f(x) = 25(4)^{\frac x 2} \]